(1–42 V1-y2–22 Compute the integral 1-y2 I-V1-22-22 (1 — x) dzdxdy
Answer #1

(1-x) dz dx dy – cycl Saya & x = st.ga - у = 2 2 2 ла- From de linites dariage - x² + y² = 1 age à 8²=1 8=1 in From x- and yUsing spesical Coordinates graph the limits in Spezica From the Coordinat es above are O limitsi - otoc 2T from Z – limits -3.5.2015, ju- TV se* (passe, game -e frangay and Zhdr = loso, 8 sin & sing, Coso) do dodo Here cc ocd oc OS RT a = 0 = b espe1-42 2T Sty ²x² (1-x) da dady = - 51-92 – 1-g2x² | [1-f sins lose] s sin & de dodø -(1-x) dz dx dy – cycl Saya & x = st.ga - у = 2 2 2 ла- From de linites dariage - x² + y² = 1 age à 8²=1 8=1 in From x- and yit = comsut - soment ut [---]n- [esu - WX9] [6502 - )